# Swarm-intelligent search for gravitational waves from eccentric binary mergers

### v3

@@ III.4 Eccentric re-searching of GWTC events
Event |
Triggered |
GPS start |
FAR |
PSO |
PyCBC |
MBTA |
GstLAL |
GW150914 |
HL |
1126259461.8 |
46.1 |
26.3 |
0.002 |
0.073 |
0.38 |
23.7 |
23.7 |
23.6 |
24.4 |
GW151012 |
HL |
1128678898.5 |
26.6 |
17.6 |
0.099 |
0.283 |
0.38 |
9.7 |
9.8 |
9.5 |
10 |
GW170104 |
HL |
1167559935.7 |
40.4 |
20.7 |
-0.034 |
0.035 |
0.38 |
13.1 |
13.2 |
13 |
13 |
GW170729 |
HL |
1185389806.8 |
70.1 |
29.1 |
0.051 |
0.325 |
2.26 |
9.5 |
9.9 |
9.8 |
10.8 |
GW170809 |
HL |
1186302518.8 |
34.7 |
27.2 |
-0.186 |
0.277 |
0.38 |
10.2 |
10.8 |
12.2 |
12.4 |
GW170814 |
HL |
1186741860.5 |
37.5 |
26.9 |
0.179 |
0.2 |
0.38 |
12.7 |
15.0 |
16.3 |
15.9 |
GW170817 |
HL |
1187008882.4 |
1.66 |
1.15 |
-0.001 |
0.01 |
0.38 |
32.4 |
32.5 |
30.9 |
33.0 |
GW170818 |
HL |
1187058326.3 |
44.8 |
22.1 |
-0.185 |
0.218 |
0.38 |
9.9 |
10.0 |
11.3 |
GW170823 |
HL |
1187529256.1 |
51.2 |
43.8 |
0.264 |
0.056 |
0.38 |
11.3 |
11.6 |
11.1 |
11.5 |
GW190403_051519 |
GW190408_181802 |
HL |
1238782699.2 |
41.9 |
17.2 |
0.002 |
0.104 |
0.38 |
14.1 |
14.1 |
13.1 |
14.4 |
14.7 |
GW190413_052954 |
HL |
1239168612.1 |
54.9 |
43.6 |
0.126 |
0.071 |
0.38 |
8.7 |
8.7 |
8.5 |
GW190413_134308 |
HL |
1239198205.2 |
82.6 |
37.6 |
-0.297 |
0 |
0.38 |
9.2 |
9.5 |
9.3 |
10.3 |
10.1 |
GW190421_213856 |
HL |
1239917954.0 |
63.5 |
40.8 |
-0.148 |
0.108 |
0.38 |
10.2 |
10.2 |
10.1 |
9.7 |
10.5 |
GW190503_185404 |
HL |
1240944861.8 |
49.4 |
34.6 |
-0.097 |
0.204 |
0.38 |
12.2 |
12.2 |
12.2 |
12.8 |
12 |
GW190512_180714 |
HL |
1241719650.7 |
27.5 |
15.9 |
-0.068 |
0.12 |
0.38 |
11.4 |
11.6 |
12.4 |
11.7 |
12.2 |
GW190513_205428 |
11.6 |
13 |
12.3 |
GW190514_065416 |
8.4 |
GW190517_055101 |
HL |
1242107479.1 |
54.4 |
24.9 |
0.365 |
0.306 |
0.38 |
9.5 |
9.5 |
10.4 |
10.8 |
11.3 |
GW190519_153544 |
HL |
1242315362.3 |
92.6 |
85.3 |
-0.499 |
0.191 |
2.63 |
11.3 |
11.3 |
13.2 |
13.7 |
12.4 |
GW190521_030229 |
HL |
1242442967.1 |
143.5 |
133.8 |
0.223 |
0.246 |
0.009 |
13.5 |
13.7 |
13.7 |
13 |
13.3 |
GW190521_074359 |
HL |
1242459857.1 |
51.8 |
43.2 |
0.162 |
0.005 |
0.38 |
24.2 |
24.4 |
24 |
22.2 |
24.4 |
GW190527_092055 |
HL |
1242984073.1 |
60.7 |
25.1 |
0.088 |
0.322 |
0.38 |
9.0 |
9.0 |
8.7 |
GW190602_175927 |
HL |
1243533584.9 |
93.1 |
72.0 |
0.036 |
0.195 |
0.38 |
12.2 |
12.2 |
11.9 |
12.6 |
12.3 |
GW190620_030421 |
10.9 |
GW190630_185205 |
15.2 |
GW190701_203306 |
HLV |
1246048404.3 |
69.9 |
56.3 |
-0.235 |
0.436 |
2.63 |
12.3 |
12.3 |
11.9 |
11.3 |
11.7 |
GW190706_222641 |
HL |
1246487219.0 |
90.3 |
48.7 |
0.148 |
0.497 |
0.38 |
13.0 |
13.0 |
11.7 |
11.9 |
12.5 |
GW190708_232457 |
13.1 |
GW190719_215514 |
HL |
1247608532.4 |
53.9 |
30.6 |
0.184 |
0.164 |
0.38 |
8.5 |
8.5 |
GW190727_060333 |
HL |
1248242631.6 |
53.3 |
48.5 |
0.151 |
0.007 |
0.38 |
11.7 |
11.7 |
11.4 |
12 |
12.1 |
GW190731_140936 |
HL |
1248617394.1 |
66.1 |
16.7 |
-0.356 |
0.09 |
0.38 |
8.6 |
8.6 |
9.1 |
8.5 |
GW190803_022701 |
HL |
1248834439.5 |
58.6 |
36.7 |
-0.119 |
0.292 |
5.64 |
8.2 |
8.5 |
8.7 |
9 |
9.1 |
GW190805_211137 |
HL |
1249074715.0 |
88.5 |
35.2 |
0.326 |
0.335 |
0.38 |
7.9 |
7.9 |
GW190828_063405 |
HL |
1251009263.1 |
39.6 |
36.1 |
0.169 |
0.164 |
0.38 |
15.7 |
15.7 |
13.9 |
15.2 |
16.3 |
GW190828_065509 |
HL |
1251010526.0 |
29.7 |
13.9 |
0.075 |
0.158 |
0.75 |
11.0 |
11.0 |
10.5 |
11.1 |
10.8 |
GW190910_112807 |
13.4 |
GW190915_235702 |
HL |
1252627040.0 |
42.2 |
28.5 |
-0.116 |
0.175 |
0.38 |
12.8 |
12.8 |
13 |
12.7 |
13 |
GW190916_200658 |
HL |
1252699636.4 |
48.9 |
35.9 |
0.05 |
0.259 |
381.12 |
6.7 |
6.8 |
8.2 |
8.2 |
GW190925_232845 |
HV |
1253489341.3 |
28.2 |
14.8 |
0 |
0.127 |
0.38 |
9.9 |
9.9 |
9 |
9.4 |
GW190926_050336 |
HL |
1253509433.6 |
87.5 |
16.2 |
-0.208 |
0.338 |
1.13 |
8.1 |
8.1 |
9 |
GW190929_012149 |
HL |
1253755327.3 |
88.3 |
46.9 |
-0.062 |
0.14 |
0.75 |
9.6 |
9.6 |
9.4 |
10.3 |
10.1 |
GW191109_010717 |
HL |
1257296855.0 |
80.4 |
59.6 |
0.176 |
0.333 |
0.38 |
16.0 |
16.0 |
13.2 |
15.2 |
15.8 |
GW191127_050227 |
HL |
1258866165.0 |
46.7 |
26.5 |
-0.214 |
0 |
1.50 |
8.7 |
8.7 |
9.5 |
9.8 |
10.3 |
GW191204_110529 |
HL |
1259492746.7 |
35.5 |
27.6 |
0.131 |
0.023 |
0.38 |
9.6 |
9.6 |
8.9 |
8.1 |
9 |
GW191215_223052 |
HL |
1260484269.4 |
39.5 |
21.1 |
0.048 |
0.085 |
0.38 |
10.6 |
10.6 |
10.3 |
10.8 |
10.9 |
GW191222_033537 |
HL |
1261020954.9 |
68.5 |
51.6 |
0.018 |
0.092 |
0.38 |
11.4 |
11.8 |
11.5 |
10.8 |
12 |
GW191230_180458 |
HL |
1261764316.2 |
90.4 |
37.9 |
0.075 |
0.109 |
0.38 |
9.8 |
9.8 |
9.6 |
9.8 |
10.3 |
GW200112_155838 |
17.6 |
GW200128_022011 |
HL |
1264213229.6 |
70.6 |
52.2 |
0.25 |
0.078 |
0.38 |
10.1 |
10.1 |
9.8 |
9.4 |
10.1 |
GW200129_065458 |
HLV |
1264316115.7 |
40.4 |
31.4 |
0.156 |
0.201 |
0.38 |
20.8 |
25.5 |
16.3 |
26.5 |
GW200208_130117 |
HLV |
1265202095.5 |
51.7 |
33.5 |
-0.13 |
0.036 |
0.38 |
10.8 |
10.8 |
9.6 |
10.4 |
10.7 |
GW200208_222617 |
8.9 |
8.2 |
GW200209_085452 |
HL |
1265273709.9 |
67.8 |
37.1 |
-0.137 |
0.149 |
0.38 |
9.0 |
9.0 |
9.2 |
9.7 |
10 |
GW200216_220804 |
HL |
1265926102.6 |
79.5 |
32.7 |
0.107 |
0.12 |
0.38 |
8.8 |
8.8 |
9 |
8.8 |
9.4 |
GW200219_094415 |
HL |
1266140671.8 |
63.7 |
39.0 |
-0.042 |
0 |
1.13 |
10.0 |
10.0 |
9.9 |
10.6 |
10.7 |
GW200220_061928 |
GW200220_124850 |
8.2 |
8.2 |
GW200224_222234 |
HLV |
1266618171.9 |
49.6 |
36.8 |
0.105 |
0.169 |
0.38 |
14.8 |
18.1 |
19.2 |
19 |
18.9 |
GW200225_060421 |
HL |
1266645877.7 |
23.0 |
18.7 |
-0.037 |
0.088 |
0.38 |
11.1 |
12.3 |
12.3 |
12.5 |
12.9 |
GW200302_015811 |
10.6 |
GW200306_093714 |
7.8 |
8.5 |
GW200308_173609 |
7.9 |
8.3 |
8.1 |
GW200311_115853 |
HLV |
1267963150.8 |
40.7 |
32.2 |
-0.019 |
0.046 |
0.38 |
15.8 |
17.1 |
17 |
16.5 |
17.7 |
GW200322_091133 |
8 |
9 |
In this section, we re-analyze the GWTC BBH events with component masses lying between 10 and 100 available in the GWOSC (Abbott et al., 2023, 2021e). A general idea behind conducting such a re-analysis is to test whether the eccentric search can produce candidates in agreement with the earlier searches. This would enable us to test our capabilities for possibly contributing to conducting follow-up studies. We find that the eccentric search recovers a majority of the confident events with either three or a two detector coincidence without much fine-tuning. Also, note that however, the false alarm rate estimates here are limited by the duration of the analyzed data over which backgrounds are estimated. The backgrounds evaluated over longer durations of data would result into better FAR estimates in general.
To illustrate this, in the case of GW190521, we further extend the search space for the component masses to 150 and use past 3.5 hours of data to compute the false alarm rate as shown in the Fig. 12. To tackle a larger search parameter space coverage, the swarm size is increased to about 2400 particles while the number of steps are kept fixed. The resulting search recovers the event with a false alarm rate close to 1 per century.
Note that we have not considered any single detector events in our analyses, which we can address in the future. In table 1, the GPS start times correspond to the average times-of-arrival across the detectors. Unlike the merger time, like in the case of GW170817 where we have used a frequency domain template, which is fixed for an event, GPS start time provided here is the start time of the triggering template and can vary depending on the duration of the g-best template corresponding to the low frequency cutoff.
@@ III.5 Runtime
Searches involving at least one neutron-star component are time-critical. To expedite the electromagnetic follow-up efforts, the LVK searches currently take a couple of seconds to few minutes to identify the event candidates. In the case of the template bank based searches, the process of matched-filtering can entirely proceed in parallel and the waveforms can be pre-computed and stored in the memory (Nitz et al., 2018). However, the speed of the PSO based searches is limited by the cost of computing templates on the fly and the serial iterations. Thus, we try to minimize the time spent at each iteration with task parallelism. We split the set of CPU intensive computations in a given iteration across the available CPU cores. Each share of the total computation is tackled by an independent process. The processes exchange the required information for computing the fitness and for communicating the results using Message Passing Interface (MPI). In other words, every core processes a fixed number of particles for all the interations. Each iteration is synchronized and a master process coordinates and drives the algorithm. The computations such as storing triggers to the disk for/or post-processing are also performed by the master process. Online analyses require that data be processed at least at the same rate as are recorded by the detectors. We have broadly timed the BNS search with 15 steps and about 5000 particles for the HL and the HLV networks as shown in Fig. 13. We observe that the runtime is almost inversely proportional to the number of CPU cores available.

### v5

@@ III.4 Eccentric re-searching of GWTC events
In this section, we re-analyze the GWTC BBH events with component masses lying between 10 and 100 available in the GWOSC (Abbott et al., 2023, 2021e). A general idea behind conducting such a re-analysis is to test whether the eccentric search can produce candidates in agreement with the earlier searches. This would enable us to test our capabilities for possibly contributing to conducting follow-up studies. We find that the eccentric search recovers a majority of the confident events with either three or a two detector coincidence without much fine-tuning. The FAR estimates here are limited by the duration ( 4096 seconds) of the data analyzed over which the backgrounds are estimated. The backgrounds evaluated over longer durations of data are expected to provide more precise FAR estimates in general.
To illustrate this, in the case of GW190521, we further extend the search space for the component masses to 150 and use past 3.5 hours of data to compute the false alarm rate as shown in Fig. 13. To tackle a larger search parameter space coverage, the swarm size is increased to about 2400 particles while the number of steps are kept fixed. The resulting search recovers the event with a FAR close to 1 per century.
@@ III.5 Runtime
Searches involving at least one neutron-star component are time-critical. To expedite the electromagnetic follow-up efforts, the LVK searches currently take a couple of seconds to few minutes to identify the event candidates. In the case of the template bank based searches, the process of matched-filtering can entirely proceed in parallel and the waveforms can be pre-computed and stored in the memory (Nitz et al., 2018). However, the speed of the PSO based searches is limited by the cost of computing templates on the fly and the serial iterations. Thus, we try to minimize the time spent at each iteration with task parallelism. We split the set of CPU intensive computations in a given iteration across the available CPU cores. Each share of the total computation is tackled by an independent process. The processes exchange the required information for computing the fitness and for communicating the results using Message Passing Interface (MPI). In other words, every core processes a fixed number of particles for all the iterations. Each iteration is synchronized and a master process coordinates and drives the algorithm. The computations such as storing triggers to the disk for/or post-processing are also performed by the master process. Online analyses require that data be processed at least at the same rate as are recorded by the detectors. We have broadly timed the BNS search with 15 steps and about 5000 particles for the HL and the HLV networks as shown in Fig. 14. We observe that the runtime is almost inversely proportional to the number of CPU cores available.